# Solar System Constraints on Light Propagation from Higher Derivative Corrections to General Relativity and the Weak Gravity Conjecture

### v1

@@ Abstract
While the two derivative action of gravitation is specified uniquely, higher derivative operators are also allowed with coefficients that are not specified uniquely by effective field theory. We focus on a four derivative operator in which the Riemann tensor couples directly to the electromagnetic field $a\,R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$. We compute the corresponding corrections to the Shapiro time delay in the solar system and compare this to data from the Cassini probe. We place an observational upper bound on the coefficient $a$ at 95% confidence $|a|<26\,(1000\,\mbox{km})^{2}$. We compare this to the weak gravity conjecture (WGC) prediction of a bound on the coefficients $a,\,b$ of four derivative operators involving the graviton and the photon; this includes the above term $a\,R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$ as well as $b\,F^{4}$. We show that by using the observed value of the $b$ coefficient from measurements of light by light scattering, which arises in the Standard Model from integrating out the electron, the WGC predicted bound for $a$ is $a\lesssim 7.8\,(1000\,\mbox{km})^{2}$. This is consistent with the above observational bound, but is intriguingly close and can be further probed in other observations.
@@ III Measurements and
where $m_{e}\approx 0.5$ MeV is the electron mass and $\alpha\approx 1/137$ is the fine structure constant. The $\sim(F^{2})^{2}$ and $\sim F^{4}$ interactions have been detected by measurements of light by light scattering at the LHC ATLAS:2017fur , with consistency with the values generated by charged particles in the Standard Model. So for them we can safely ignore any $b_{\mbox{\tiny{UV}}}$ contributions at low energies. So we have
@@
Integrating out the electron also generates the $\sim R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta},\,R_{\mu\nu}F^{\mu\alpha}F^{\nu}_{\,\,\alpha},\,RF^{2}$ terms with shifted values $\Delta a,\Delta a^{\prime},\,\Delta a^{\prime\prime}\sim\alpha/m_{e}^{2}$ Drummond:1979pp . If this is the correct value for $a,\,a^{\prime},\,a^{\prime\prime}$, then it is easy to check that their effects are so small in the solar system and are essentially unmeasurable. However, since we currently do not have a measurement of these coefficients, we do not know that this is the correct value. All we know is that the value of $a,\,a^{\prime},\,a^{\prime\prime}$ will change, or “flow”, by $\Delta a,\,\Delta a^{\prime},\,\Delta a^{\prime\prime}$ above and below the electron mass. It could be that their actual values are much larger than this (positive or negative), and the correction from integrating out the electron is a negligible correction. So for the purpose of this paper, we shall consider the possibility that $a,\,a^{\prime},\,a^{\prime\prime}$ are large as provided by their UV values. So we shall take
@@
so that $M\leq\sqrt{2}\,|Q|M_{\mbox{\tiny{Pl}}}$ for extremal black holes, just as elementary particles should obey $m\leq\sqrt{2}\,e\,M_{\mbox{\tiny{Pl}}}$ in the original form of the WGC.
@@
In the approximate vacuum of the solar system away from the sun, we have $R_{\mu\nu}\approx 0$; so the $a^{\prime}$ term will not play a significant role. While we still have a non-zero $R_{\mu\nu\alpha\beta}$ from the sun; so the $a$ term will be essential. The above WGC inequality only refers to the sum, but barring some highly unexpected cancellation [^1] For example, integrating out the electron at 1-loop leads to $\Delta a^{\prime}=-13\Delta a$ Drummond:1979pp . Also, the coefficients should obey $a^{\prime}=-4a$ to avoid higher derivative equations of motion. So in all known examples to us, there is no cancellation of the form $a=-a^{\prime}$., the WGC bound can be interpreted roughly as $a\lesssim 7.8(1000\,\mbox{km})^{2}$. So we will be effectively referring to just $a$ from now on.
In this work we will wonder what happens if nature saturates or is close to this bound, i.e., $a\sim 7.8(1000\,\mbox{km})^{2}$. While we do not have a good a priori reason to expect nature to be near the bound, it does provide the sharpest test of the WGC. We shall check with existing data whether this is compatible with current solar system tests of gravity.
@@ IV Altered Light Travel Time
One of, if not the, most precise tests of general relativity is the prediction of an increase in the round trip time of light due to a massive object, such as the sun. This is the famous Shapiro time delay Shapiro:1964uw . In the presence of the higher derivative correction $a\,R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$ this prediction will be altered.
@@
So in order to avoid superluminality, we need $16|a|/b^{2}<$ 1. As described earlier, we will be interested in values of $a$ in the vicinity of $a\sim 10\,(1000\,\mbox{km})^{2}$, while the smallest impact parameter we will study is $b_{\mbox{\tiny{min}}}=1.6\,R_{\odot}\approx 1.1\times 10^{6}\,$km. Hence we have $6|a|/b_{\mbox{\tiny{min}}}^{2}\sim 5\times 10^{-5}$, ensuring that the behavior is always subluminal (away from the maximum speed correction, the corrections from this new operator are even smaller).
@@ VII Discussion
Further work can include considering other observations to constrain the term $aR_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$. But also to consider observations to constrain $a^{\prime}R_{\mu\nu}F^{\mu\alpha}F^{\nu}_{\,\,\alpha}$. The latter requires considering regions in which there is matter, so that the Ricci tensor is non-zero. This is important since the WGC is actually sensitive to the sum $a+a^{\prime}$. Alternatively, one could try to sharpen the argument that $a^{\prime}=-4a$ to avoid ghosts in the UV, which would link the coefficients.

### v2

# Solar System Constraints on Light Propagation from Higher Derivative Corrections to General Relativity and Implications for Quantum Gravity

@@ Abstract
While the two derivative action of gravitation is specified uniquely, higher derivative operators are also allowed with coefficients that are not specified uniquely by effective field theory. We focus on a four derivative operator in which the Riemann tensor couples directly to the electromagnetic field $a\,R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$. We compute the corresponding corrections to the Shapiro time delay in the solar system and compare this to data from the Cassini probe. We place an observational upper bound on the coefficient $a$ at 95% confidence $|a|<26\,(1000\,\mbox{km})^{2}$. By way of motivation, we also compare this to a weak gravity conjecture (WGC) prediction of a bound on the coefficients $a,\,b$ of four derivative operators involving the graviton and the photon; this includes the above term $a\,R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$ as well as $b\,F^{4}$. We show that by using the observed value of the $b$ coefficient from measurements of light by light scattering, which arises in the Standard Model from integrating out the electron, the WGC predicted bound for $a$ is $a\lesssim 7.8\,(1000\,\mbox{km})^{2}$. This is consistent with the above observational bound, but is intriguingly close and can be further probed in other observations.
@@ III Measurements and
where $m_{e}\approx 0.5$ MeV is the electron mass and $\alpha\approx 1/137$ is the fine structure constant. The $\sim(F^{2})^{2}$ and $\sim F^{4}$ interactions have been detected by measurements of light by light scattering at the LHC ATLAS:2017fur ; TOTEM:2021zxa , with consistency with the values generated by charged particles in the Standard Model. So for them we can safely ignore any $b_{\mbox{\tiny{UV}}}$ contributions at low energies. So we have
@@
Integrating out the electron also generates the $\sim R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta},\,R_{\mu\nu}F^{\mu\alpha}F^{\nu}_{\,\,\alpha},\,RF^{2}$ terms with shifted values $\Delta a,\Delta a^{\prime},\,\Delta a^{\prime\prime}\sim\alpha/m_{e}^{2}$, as computed in Ref. Drummond:1979pp (updates include Ref. Bastianelli:2008cu ). If this is the correct value for $a,\,a^{\prime},\,a^{\prime\prime}$, then it is easy to check that their effects are so small in the solar system and are essentially unmeasurable. However, since we currently do not have a measurement of these coefficients, we do not know that this is the correct value. All we know is that the value of $a,\,a^{\prime},\,a^{\prime\prime}$ will change, or “flow”, by $\Delta a,\,\Delta a^{\prime},\,\Delta a^{\prime\prime}$ above and below the electron mass. It could be that their actual values are much larger than this (positive or negative), and the correction from integrating out the electron is a negligible correction. So for the purpose of this paper, we shall consider the possibility that $a,\,a^{\prime},\,a^{\prime\prime}$ are large as provided by their UV values. So we shall take
@@
so that $M\leq\sqrt{2}\,|Q|M_{\mbox{\tiny{Pl}}}$ for extremal black holes, just as elementary particles should obey $m\leq\sqrt{2}\,e\,M_{\mbox{\tiny{Pl}}}$ in the original form of the WGC. We note that other interesting developments include Refs. Cheung:2014ega ; Cremonini:2009ih ; Bellazzini:2019xts ; Cremonini:2019wdk ; Noumi:2022ybv ; BarrosoVarela:2023ull ; Abe:2023anf ; Bittar:2024xuc ; Barbosa:2025uau .
@@
In the approximate vacuum of the solar system away from the sun, we have $R_{\mu\nu}\approx 0$; so the $a^{\prime}$ term will not play a significant role. While we still have a non-zero $R_{\mu\nu\alpha\beta}$ from the sun; so the $a$ term will be essential. The above WGC inequality only refers to the sum, but barring some highly unexpected cancellation [^1] For example, integrating out the electron at 1-loop leads to $\Delta a^{\prime}=-13\Delta a$ Drummond:1979pp . Also, the coefficients should obey $a^{\prime}=-4a$ to avoid higher derivative equations of motion. Or we should have $a^{\prime}=-2a$ to form the Weyl tensor. So in all known examples to us, there is no cancellation of the form $a=-a^{\prime}$., the WGC bound can be interpreted roughly as $a\lesssim 7.8(1000\,\mbox{km})^{2}$. So we will be effectively referring to just $a$ from now on. In fact one can go further (see Refs. Hamada:2018dde ; Alberte:2020bdz ): by performing some field redefinitions, one can take all the $a,\,a^{\prime},\,a^{\prime\prime}$ terms to be combined into the form $-a\,W_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$, where $W_{\mu\nu\alpha\beta}$ is the Weyl tensor (giving $a^{\prime}=-2a$ and $a^{\prime\prime}=a/3$); this makes this a more direct photon-graviton coupling. (Terms of the form $R_{\mu\nu}F^{\mu\alpha}F^{\nu}_{\,\,\alpha}$ and $RF^{2}$ are not true graviton-photon couplings in a sense: This is because the Einstein equations tell us to leading order $R_{\mu\nu}$ is determined by $T_{\mu\nu}$, which in turn is determined by a bilinear in $F_{\mu\nu}$ the electromagnetic field strength in vacuum. Upon substitution this means the $a^{\prime},\,a^{\prime\prime}$ terms, to leading order, play a redundant role to the terms quartic in $F_{\mu\nu}$, which we are already including.) Alternatively, one can set $a^{\prime}=-4a$ and $a^{\prime\prime}=a$ to achieve two derivative equations of motion Horndeski:1976gi ; Barrow:2012ay ; BeltranJimenez:2013btb . In either way, the values of $a^{\prime},\,a^{\prime\prime}$ terms would be uniquely determined by $a$.
In this work, we are generically interested in placing observational bounds on these corrections. And, by way of motivation, we will wonder what happens if nature saturates or is close to this theoretical bound, i.e., $a\sim 7.8(1000\,\mbox{km})^{2}$. While we do not have a good a priori reason to expect nature to be near this theoretical bound, it does provide a test of the WGC. We shall check with existing data whether this is compatible with current solar system tests of gravity.
@@ IV Altered Light Travel Time
One of, if not the, most precise tests of general relativity is the prediction of an increase in the round trip time of light due to a massive object, such as the sun. This is the famous Shapiro time delay Shapiro:1964uw . In the presence of the higher derivative correction $a\,R_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$ this prediction will be altered. Interesting other works exploring the observational consequences of this operator includes Refs. Chen:2016hil ; Li:2017zsb ; Li:2018jjk ; Abbas:2021whh ; Zhang:2021hit ; Jana:2021lqe ; Bamba:2008ja .
@@
So in order to avoid superluminality, we need $16|a|/b^{2}<$ 1. As described earlier, we will be interested in values of $a$ in the vicinity of $a\sim 10\,(1000\,\mbox{km})^{2}$, while the smallest impact parameter we will study is $b_{\mbox{\tiny{min}}}=1.6\,R_{\odot}\approx 1.1\times 10^{6}\,$km. Hence we have $6|a|/b_{\mbox{\tiny{min}}}^{2}\sim 5\times 10^{-5}$, ensuring that the behavior is always subluminal (away from the maximum speed correction, the corrections from this new operator are even smaller). We note that other discussions of superluminality in related contexts, includes Refs. Daniels:1993yi ; Dittrich:1998fy ; Novello:1999pg ; Shore:2007um ; Hollowood:2008kq ; Hollowood:2015elj ; Goon:2016une ; Hertzberg:2017abn ; deRham:2020zyh .
@@ VII Discussion
Further work can include considering other observations to constrain the term $aR_{\mu\nu\alpha\beta}F^{\mu\nu}F^{\alpha\beta}$. But also to consider observations to constrain $a^{\prime}R_{\mu\nu}F^{\mu\alpha}F^{\nu}_{\,\,\alpha}$. The latter requires considering regions in which there is matter, so that the Ricci tensor is non-zero. This is important since the WGC is actually sensitive to the sum $a+a^{\prime}$. Alternatively, one could try to sharpen the argument that $a^{\prime}=-4a$ to avoid ghosts in the UV, which would link the coefficients. Or use the Weyl tensor coupling, which links the coefficients as $a^{\prime}=-2a$, and provides a more direct probe of graviton-photon coupling.